OpenStudy (samara8954):

The value of a professional basketball player's autograph rose 30% in the last year. It is now worth $273.00. What was it worth a year ago? $200.00 $210.00 $220.00 $230.00

OpenStudy (samara8954):

@AlexandervonHumboldt2

jimthompson5910 (jim_thompson5910):

x = last year's value (last year's value) + (30% of last year's value) = this year's value x + 30% of x = 273 x + 0.30x = 273 solve for x

OpenStudy (alexandervonhumboldt2):

x*(1+0.30)=273

OpenStudy (danjs):

\[273 = \frac{ (100 + 30) *x}{ 100 }\]

OpenStudy (alexandervonhumboldt2):

210